# Linear Discriminant Regularized Regression

### v2

@@ 1 Introduction
For any numbers $a,b\in\mathbb{R}$, we write $a\vee b=\max\{a,b\}$. For any vector $v$ and $1\leq q\leq\infty$, we use $\|v\|_{q}$ to denote the standard $\ell_{q}$-norm. For any positive integer $d$, we write $[d]:=\{1,\ldots,d\}$. For any matrix $Q\in\mathbb{R}^{d_{1}\times d_{2}}$, any $i\in[d_{1}]$ and $j\in[d_{2}]$, we write $Q_{i\cdot}$ for its $i$th row and $Q_{j}$ for its $j$th column. For norms, we write $\|Q\|_{\infty}=\max_{i,j}|Q_{ij}|$, $\|Q\|_{F}^{2}=\sum_{i,j}Q_{ij}^{2}$, $\|Q\|_{1}=\sum_{i,j}|Q_{ij}|$, $\|Q\|_{1,2}=\sum_{i}\|Q_{i\cdot}\|_{2}$ and $\|Q\|_{\rm op}=\sup_{v:\|v\|_{2}=1}\|Qv\|_{2}$. For any symmetric, semi-positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we use $\lambda_{1}(Q),\lambda_{2}(Q),\ldots,\lambda_{d}(Q)$ to denote its eigenvalues in non-increasing order. For any sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if there exists some constant $C>0$ such that $a_{n}\leq Cb_{n}$, and $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
@@ 2 Methodology
We discuss our classification procedure in this section. Our goal is to estimate the Bayes rule in (1.2) which, for a given $x\in\mathbb{R}^{p}$, requires estimation of the discriminant functions
$G_{\ell}(x):=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2x^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}),\qquad\text{for all }\ell\in[L].$ (2.1)
This, in turn, requires to estimate the discriminant directional matrix $B^{*}=\Sigma_{w}^{-1}{M}$. The following lemma establishes the connection between $B^{*}$ and the regression matrix $B$ in (1.12). Recall the definitions of $\pi$, $D_{\pi}$ and ${M}$ from (1.3) as well as $\Sigma$ from (1.4).
@@
Next, using
$\Sigma=\Sigma_{w}+\Sigma_{b}=\Sigma_{w}+{M}D_{\pi}{M}^{\top},$ (2.6)
@@
Lemma 1 suggests that estimation of $B^{*}$ could be done by first estimating the matrix $B$ and then estimating $H^{-1}$ according to (2.3). We discuss these two steps in details in the next two sections.
@@ 2.1 Estimation of the regression matrix
To accommodate high-dimensional data, we estimate $B$ via regularized regression methods in (1.14). Choices of the penalty term $\text{pen}(B)$ depend on the concrete problem at hand. For instance, when entries of $B$ exhibit certain smoothness structure such as in some imaging or audio applications (Hastie et al., 1995), one could choose $\text{pen}(B)={\rm tr}(B^{\top}\Omega B)$ for some pre-specified “roughness” penalty matrix $\Omega$. In this paper we mainly focus on the following two types of structural estimation of $B$.
Sparsity-based structural estimation.
@@
$\text{pen}(B)$ $=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{1}$ (2.7)
for some tuning parameter $\lambda>0$. In Section 3.3 we provide a complete analysis of this estimator. If there exist many features that are not predictive for all levels of the response label, then the group-lasso penalty $\text{pen}(B)=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ (Yuan and Lin, 2006) is a more suitable choice. In this case, one can retain a much smaller subset of important features and achieve feature selection. Furthermore, one can also consider either a row-wise or a column-wise elastic net penalty (Zou and Hastie, 2005) if the features tend to (highly) correlate with each other. Our numerical results in Sections 4 and 5 include both the lasso penalty and the elastic net penalty.
Low-rank structural estimation.
@@
We conclude this section by giving another concrete example under which $B$ has low rank. In Bing and Wegkamp (2023) the authors consider the model $X=AZ+W$ with latent factors $Z\in\mathbb{R}^{r}$ for some $r\ll p$ and some general additive noise $W$, that is independent of both $Z$ and the label $Y$. Conditioning on $Y={\bm{e}}_{\ell}$ for all $\ell\in[L]$, the low-dimensional factors $Z$ are assumed to follow Gaussian distributions with different conditional means $\alpha_{\ell}\in\mathbb{R}^{r}$ but the same within-class covariance matrix $\Sigma_{z}\in\mathbb{R}^{r\times r}$. When $W$ is also Gaussian, this model is a sub-model of (1.1) with $\mu_{\ell}=A\alpha_{\ell}$ and $\Sigma_{w}=A\Sigma_{z}A^{\top}+{\rm Cov}(W)$. When $r\leq\min\{p,L\}$, it is easy to verify that ${\rm rank}(B)\leq r$. The two-step approach in Bing and Wegkamp (2023) uses the unsupervised Principal Component Analysis (PCA) to reduce the feature dimension and then base classification on the reduced dimension. It is entirely different from our approach in Section 2.2 but connected to the one in Section 2.3 in the way that our approach in Section 2.3 is also a two-step procedure but reduces the feature space in the first step in a supervised way via the regression step in (1.14). Numerical comparison of both procedures are also included in our simulation study in Appendix A.
2.2 Estimation of the discriminant matrix and the Bayes rule
The discriminant functions in (2.1) that form the basis of the Bayes classification rule (1.2) necessitate that we estimate the matrix of the conditional means ${M}$, the vector of prior probabilities $\pi$ as well as the discriminant directional matrix $B^{*}$. We adopt the following standard estimators
@@ 2.2 Estimation of the discriminant matrix and the Bayes rule
to estimate ${M}$ and $D_{\pi}$, respectively. Recall from Lemma 1 that $B^{*}=BH^{-1}$. Since the estimator $\widehat{B}$ of $B$ is computed in (1.14), in view of the closed-form of $H$ in (2.3), we propose to estimate $H$ by the plug-in estimator
$\widehat{H}=D_{\widehat{\pi}}-{1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}{\mathbf{X}}\widehat{B}={1\over n}\left({\mathbf{Y}}^{\top}{\mathbf{Y}}-\widehat{B}^{\top}{\mathbf{X}}^{\top}{\mathbf{X}}\widehat{B}\right).$ (2.10)
Note that ${\mathbf{X}}\widehat{B}$ is simply the in-sample fit of the regression in (1.14). Computation of ${\mathbf{X}}\widehat{B}$ thus can be done efficiently even in non-parametric regression settings, for instance, when the columns of ${\mathbf{X}}$ consist of basis expansions. Now, using $B^{*}=BH^{-1}$, we can further estimate $B^{*}$ by
@@
Here for any matrix $M$, $M^{+}$ denotes its moore-penrose inverse. Finally, in view of (2.1), we estimate the discriminant functions by
$\widehat{G}_{\ell}(x)=\widehat{\mu}_{\ell}^{\top}\widehat{B}^{*}_{\ell}-2x^{\top}\widehat{B}^{*}_{\ell}-2\log(\widehat{\pi}_{\ell}),\qquad\text{for all $\ell\in[L]$. }$ (2.12)
The corresponding classifier is
$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad\text{ for any $x\in\mathbb{R}^{p}$}.$ (2.13)

### v3

@@ 1.3 Notation
For any numbers $a,b\in\mathbb{R}$, we write $a\vee b=\max\{a,b\}$. For any vector $v$ and $1\leq q\leq\infty$, we use $\|v\|_{q}$ to denote the standard $\ell_{q}$-norm. For any positive integer $d$, we write $d]:=\{1,\ldots,d\}$. For any matrix $Q\in\mathbb{R}^{d_{1}\times d_{2}}$, any $i\in[d_{1}]$ and $j\in[d_{2}]$, we write $Q_{i\cdot}$ for its $i$th row and $Q_{j}$ for its $j$th column. For norms, we write $\|Q\|_{\infty}=\max_{i,j}|Q_{ij}|$, $\|Q\|_{F}^{2}=\sum_{i,j}Q_{ij}^{2}$, $\|Q\|_{1}=\sum_{i,j}|Q_{ij}|$, $\|Q\|_{1,2}=\sum_{i}\|Q_{i\cdot}\|_{2}$ and $\|Q\|_{\rm op}=\sup_{v:\|v\|_{2}=1}\|Qv\|_{2}$. For any symmetric, semi-positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we use $\lambda_{1}(Q),\lambda_{2}(Q),\ldots,\lambda_{d}(Q)$ to denote its eigenvalues in non-increasing order. For any sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if there exists some constant $C>0$ such that $a_{n}\leq Cb_{n}$, and $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
@@ 2 Methodology
We discuss our classification procedure in this section. Our goal is to estimate the rule in (1.1) which entails estimating the discriminant functions:
$\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2x^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell})=\mu_{\ell}^{\top}B^{*}_{\ell}-2x^{\top}B^{*}_{\ell}-2\log(\pi_{\ell}),\quad\text{for all }\ell\in[L].$ (2.1)
Note that there is no need to estimate the $p\times p$ matrix $\Sigma_{w}^{-1}$ explicitly. The following lemma establishes the connection between the discriminant directional matrix $B^{*}=(B_{1}^{*},\cdots,B_{L}^{*})$ in (1.7) and the regression matrix $B$ in (1.8). Write $\pi=(\pi_{1},\ldots,\pi_{L})^{\top}$ and $D_{\pi}={\rm diag}(\pi)\in\mathbb{R}^{L\times L}$.
@@
Next, using the fact that
$\Sigma=\Sigma_{w}+{M}D_{\pi}{M}^{\top},$ (2.6)
@@
Lemma 1 suggests that estimation of $B^{*}$ could be done by first estimating the matrix $B$ and then estimating $H^{-1}$ according to (2.3). We discuss these two steps in detail in the next two sections.
@@ 2.1 Estimation of the regression matrix
To accommodate high-dimensional data, we estimate $B$ via penalized regression methods in (1.10). Choices of the penalty term depend on the concrete problem at hand. For instance, when entries of $B$ exhibit certain smoothness structure such as in some imaging or audio applications (Hastie et al., 1995), one could choose $\text{pen}(B)={\rm tr}(B^{\top}\Omega B)$ for some pre-specified “roughness” penalty matrix $\Omega$. In this paper we mainly focus on the following two types of structural estimation of $B$.
2.1.1 Sparsity-based structural estimation.
@@
$\text{pen}(B)$ $=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{1}=\lambda\sum_{\ell=1}^{L}\|B_{\ell}\|_{1}$ (2.7)
for some tuning parameter $\lambda>0$. In Section 3.3 we provide a complete analysis of this estimator. If there exist many features that are not predictive for all levels of the response label, then the group-lasso penalty $\text{pen}(B)=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ (Yuan and Lin, 2006) is a more suitable choice. In this case, one can retain a much smaller subset of important features and achieve feature selection. From Lemma 1 we see that row-sparsity of $B$ is equivalent to that of $B^{*}$. Our focus in this paper, however, is on prediction, not feature selection. Furthermore, one can also consider either a row-wise or a column-wise elastic net penalty (Zou and Hastie, 2005) if the features tend to (highly) correlate with each other. Our numerical results in Sections 4 and 5 include both the lasso penalty and the elastic net penalty.
2.1.2 Low-rank structural estimation.
@@
We conclude this section by giving another concrete example under which $B$ has low rank. In Bing and Wegkamp (2023) the authors consider the model $X=AZ+W$ with latent factors $Z\in\mathbb{R}^{r}$ for some $r\ll p$ and some general additive noise $W$, that is independent of both $Z$ and the label $Y$. Conditioning on $Y={\bm{e}}_{\ell}$ for all $\ell\in[L]$, the low-dimensional factors $Z$ are assumed to follow Gaussian distributions with different conditional means $\alpha_{\ell}\in\mathbb{R}^{r}$ but the same within-class covariance matrix $\Sigma_{z}\in\mathbb{R}^{r\times r}$. When $W$ is also Gaussian, this model is a sub-model of (3.1) with $\mu_{\ell}=A\alpha_{\ell}$ and $\Sigma_{w}=A\Sigma_{z}A^{\top}+{\rm Cov}(W)$. When $r\leq\min\{p,L\}$, it is easy to verify that ${\rm rank}(B)\leq r$. The two-step approach in Bing and Wegkamp (2023) uses the unsupervised Principal Component Analysis (PCA) to reduce the feature dimension and then base classification on the reduced dimension. It is entirely different from our approach in Section 2.2 but connected to the one in Section 4.1 in the way that our approach in Section 4.1 is also a two-step procedure but reduces the feature space in the first step in a supervised way via the regression step in (1.10). Numerical comparison of both procedures are also included in our simulation study in Section 4.3.2.
2.2 Estimation of the discriminant functions
The discriminant functions in (2.1) necessitates that we estimate the conditional mean matrix M{M} and the discriminant directional matrix $B^{*}$. The latter also requires estimating the vector of prior probabilities $\pi$. We adopt the following standard estimators
@@ 2.2 Estimation of the discriminant functions
to estimate M{M} and $D_{\pi}$, respectively. Recall from Lemma 1 that $B^{*}=BH^{-1}$. Since the estimator $\widehat{B}$ of $B$ is computed in (1.10), in view of the closed-form of $H$ in (2.3), we propose to estimate $H$ by the plug-in estimator
[…]
Note that $^{\mathbf{X}}\widehat{B}$ is simply the in-sample fit of the regression in (1.10). Computation of $^{\mathbf{X}}\widehat{B}$ thus can be done efficiently even in non-parametric regression settings, for instance, when the columns of $\mathbf{X}$} consist of basis expansions. Now, using $B^{*}=BH^{-1}$, we can further estimate $B^{*}$ by
@@
Here for any matrix $M$, $M^{+}$ denotes its Moore-Penrose inverse. Finally, we estimate the discriminant functions in (2.1) by
$\widehat{\mu}_{\ell}^{\top}\widehat{B}^{*}_{\ell}-2x^{\top}\widehat{B}^{*}_{\ell}-2\log(\widehat{\pi}_{\ell}),\qquad\text{for all $\ell\in[L]$,}$ (2.12)